OpenStudy (mathmath333):

quickest way to find remainder for \(\LARGE \begin{align} \color{black}{ \dfrac{65^{{37}^{80}}}{39} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (ikram002p):

waitiiiing :D

OpenStudy (ikram002p):

are you gonna show answer or should i try ?

OpenStudy (mathmath333):

my method was long , so i want quickest (lazy) way

OpenStudy (ikram002p):

\(\Large \begin{align*} 65 &\equiv 26 \mod 39 \\ 65^{37^{80}} &\equiv 26 ^{37^{80}}\mod 39 \\ \text{now note this } \\ 26^{\phi(39)} &\equiv 1 \mod 39\\ 26^{24}&\equiv 1 \mod 39 \\ 26^{24 n}&\equiv 1^n \mod 39 \\ 26^{24 n+q}&\equiv 1\times 26^q \mod 39 \\ 26^{24 n+q}&\equiv 26^q \mod 39 \\ \text { so we need to know q}\\ 37^{80} \mod 24 &\equiv q \mod 24 \\ 37^{80}&\equiv 13^{80} \mod 24 \\ 13^{\phi(24)}&\equiv 1 \mod 24 \\ 13^{8}&\equiv 1 \mod 24 \\ 13^{80}&\equiv 1 \mod 24 \\ \text{thus } q=1 \\ \end{align*} \)

OpenStudy (ikram002p):

by understanding that relation then you can quick reach this result :- \(\Large \begin{align*} 65^{37^{80}} &\equiv 26^{37^{80}}\mod 39\\ &\equiv 26^{({37^{80}\mod 24})}\mod 39 \\ &\equiv 26^{1}\mod 39\\ &\equiv 26 \mod 39 \end{align*} \)

OpenStudy (mathmath333):

\(\Large \begin{align} \color{black}{\normalsize \text{chinese remainder theorm}\hspace{.33em}\\~\\ \dfrac{65^{{37}^{80}}}{39} \hspace{.33em}\\~\\ \dfrac{65^{{37}^{80}}}{39}=3r_2x+13r_1y \hspace{.33em}\\~\\ 3x+13y=1\hspace{.33em}\\~\\ x=-4,y=1\hspace{.33em}\\~\\ \dfrac{65^{{37}^{80}}}{39}=-12r_2+13r_1 \hspace{.33em}\\~\\ 39=13\times 3 \hspace{.33em}\\~\\ \dfrac{65^{{37}^{80}}}{3}=r_1 \hspace{.33em}\\~\\ =\dfrac{(3\times 22-1)^{{37}^{80}}}{3}\hspace{.33em}\\~\\ =\dfrac{(-1)^{{37}^{80}}}{3}\hspace{.33em}\\~\\ =\dfrac{(-1)^{odd}}{3}\hspace{.33em}\\~\\ =\dfrac{-1}{3}\hspace{.33em}\\~\\ =\dfrac{2}{3}\hspace{.33em}\\~\\ =2\hspace{.33em}\\~\\ \fbox{r_1=2 }\hspace{.33em}\\~\\ \dfrac{65^{{37}^{80}}}{13}=r_2 \hspace{.33em}\\~\\ =\dfrac{(13\times 5)^{{37}^{80}}}{13}\hspace{.33em}\\~\\ =\dfrac{(0)^{{37}^{80}}}{13}\hspace{.33em}\\~\\ =0\hspace{.33em}\\~\\ \fbox{r_2=0 }\hspace{.33em}\\~\\ \dfrac{65^{{37}^{80}}}{39}=-12(0)+13(2)=\LARGE 26 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (ikram002p):

hmm interesting :)

OpenStudy (mathmath333):

Yes but it seems little longer...

OpenStudy (mathmath333):

it works in case the divisors are composite

OpenStudy (ikram002p):

meybe you wanna have a look @ganeshie8

ganeshie8 (ganeshie8):

ikram's method looks more simpler..

ganeshie8 (ganeshie8):

chinese remainder thm looks fancy and interesting xD

OpenStudy (mathmath333):

yes but its still like bolt from the blue for me

OpenStudy (ikram002p):

he asked for quick :O

ganeshie8 (ganeshie8):

she is using only one thm : euler generalizatio to fermat's little thm

OpenStudy (ikram002p):

i can solve using Eclids algorithm without caring how much its gonna take time

ganeshie8 (ganeshie8):

you must be knowing fermat's little thm for prime moduli : \[a^{p-1}\equiv 1\pmod{p}\] ?

OpenStudy (mathmath333):

yes it quick compared to the chineese.

OpenStudy (mathmath333):

yes

ganeshie8 (ganeshie8):

Notice fermat's little thm only works for prime moduli euler's generalization is a nice extension to fermat's little thm

ganeshie8 (ganeshie8):

eulero's generalization : \[a^{\phi(n)}\equiv 1\pmod{n}\] here \(n\) could be any integer

ganeshie8 (ganeshie8):

(any \(n\) relatively prime to \(a\) )

OpenStudy (ikram002p):

hehe exactly :P

OpenStudy (mathmath333):

yes i learnt that \(\phi(n)=n(1-1/p_1)(1-1/p_1).....(1-1/p_n)\) where p1,p2...pn are prime factors

ganeshie8 (ganeshie8):

\(\phi(n)\) is a very friendly function which you can compute easily w/o any formulas

ganeshie8 (ganeshie8):

\(\phi(n)\) gives the number of positive integers that are less than \(n\) and relatively prime to \(n\)

OpenStudy (mathmath333):

for euler method a and n must be co prime ?

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

lets do an example of phi function maybe ?

OpenStudy (mathmath333):

yes

ganeshie8 (ganeshie8):

\(\phi(6) = 2\) because there are exactly \(2\) positive integers less than \(6\) that are relatively prime to \(6\) : \( \{1, 5\}\)

ganeshie8 (ganeshie8):

\(\phi(7) = ?\)

OpenStudy (mathmath333):

6

ganeshie8 (ganeshie8):

because every positive integer less tha a prime is relatively prime

OpenStudy (mathmath333):

7(1-1/7)=6

ganeshie8 (ganeshie8):

that formula works notice also \(\phi(p) = p-1\)

OpenStudy (mathmath333):

its for prime \(p\)

ganeshie8 (ganeshie8):

yes

OpenStudy (mathmath333):

then it becomes fermat

ganeshie8 (ganeshie8):

Exactly! lets see if we can work below problem using euler thm : find the remainder when \(35^{401}\) is divided by \(12\)

ganeshie8 (ganeshie8):

in other words, reduce below : \[x\equiv 35^{401} \pmod{12}\]

OpenStudy (mathmath333):

\(\dfrac{(-1)^{401}}{12}=11\)

ganeshie8 (ganeshie8):

Very smart :) thats a bad example haha let me cook a harder one lol

ganeshie8 (ganeshie8):

\[x\equiv 41^{401} \pmod{12}\]

OpenStudy (mathmath333):

w8

ganeshie8 (ganeshie8):

kk try to use euler thm

OpenStudy (mathmath333):

\(\Large \dfrac{41^{401}}{12}\\ =\dfrac{5^{401}}{12}\\ =\dfrac{5\cdot 5^{400}}{12}\\ =\dfrac{5\cdot 25^{200}}{12}\\ =\dfrac{5\cdot 1^{200}}{12}\\ =5 \)

OpenStudy (mathmath333):

oh euler

ganeshie8 (ganeshie8):

thats clever lol

ganeshie8 (ganeshie8):

euler is not needed for working that but you may use it just to get some practice

OpenStudy (mathmath333):

i was having difficulty for this type \(\Large a^{b^c}\)

OpenStudy (mathmath333):

may be i need more practice with euler. thnks

ganeshie8 (ganeshie8):

\(\phi(12) = 4\) so euler gives us \(41^{4}\equiv 1 \pmod{12}\) thats the end of euler, we simply use that to reduce the power

ganeshie8 (ganeshie8):

\[41^{401} \equiv 41(41^{400}) \equiv 41(41^4)^{100}\equiv 41(1)^{100}\equiv 41 \equiv 5 \pmod{12}\]